How do you calculate the load and bleeder currents in a voltage divider? Vs = 10V. R1 & R2 both 10,000 ohms each.

3. Connect a 100 kS2 load to the circuit. Apply power, and measu age (½). Use la to calculate: . The load and bleeder currents.

media%2F64a%2F64a2e965-add2-471e-b520-65

media%2Fc73%2Fc7338f30-aedc-484c-ad5a-98
Answer #1

Ans) In the above given circuit ,it is given that R_{1}=R_{2}=10000\: \: \Omega

and R_{L}=100\: \: k\Omega

Using voltage division rule ,the load voltage in the above circuit is given as

V_{L}=V_{s}*\frac{R_{2}\parallel R_{L}}{R_{1}+R_{2}\parallel R_{L}}

V_{L}=V_{s}*\frac{\frac{R_{2}* R_{L}}{R_{2}+ R_{L}}}{R_{1}+\frac{R_{2}* R_{L}}{R_{2}+ R_{L}}}

V_{L}=V_{s}*\frac{R_{2}* R_{L}}{R_{1}(R_{2}+ R_{L})+R_{2}* R_{L}}

Now substitute the values we get

V_{L}=10*\frac{10000* 100k}{10000*(10000+ 100k)+10000*100k}

V_{L}=4.762\: \: V

The bleeder current is given as

I_{Bleeder}=\frac{V_{L}}{R_{2}}=\frac{4.762}{10000}=476.2\: \: \mu A

I_{Bleeder}=476.2\: \: \mu A

Load current is given as

I_{L}=\frac{V_{L}}{R_{L}}=\frac{4.762}{100k}=47.62\: \: \mu A

I_{L}=47.62\: \: \mu A